The partial pressures of the gases in a 22.414 L volume of air at STP are: oxygen, 21.224 kPa; nitrogen, 79.119 kPa; argon 0.946 kPa; and carbon dioxide, 0.036 kPa. Determine the gram-molecular weight of air


1
Expert's answer
2021-06-07T03:15:07-0400

Total pressure= pO2+Pn2+pAr+Pco2

Total pressure= 21.2+79.1+0.94+0.036

Total pressure= 101.27kpa=0.99

PV=nRT

At STP,T= 298k

0.99×22.4=n×1/12×298

n= 0.89